Two isolated conducting spheres are separated by a large distance. Sphere 1 has a radius of 20 cm and an initial charge 30 nC while Sphere 2 has a radius of 60 cm and an initial charge 70 nC. A very thin copper wire is now connected to the spheres to allow charge to flow between them. How much charge will be transferred from Sphere 2 to Sphere 1? (Note that the charge transferred can be positive, negative or zero.)
Answer #1

Dear student,

Find this solution, and RATE IT ,If you find it is helpful .your rating is very important to me.If any incorrectness ,kindly let me know I will rectify them soon.

Thanks for asking ..Ri=200m 9=300C R2 = 60cm, 2 = 7016 YA = VB 3 Kw - K ) 2 = 60 a aloo, T +/2 = 4 +22= 100 nC from ② f1 9 + 39 = loone ④ inc Cha